# Radial Gradients and Intrinsic Scatter in MaNGA Galaxies

### v1

@@ 5.2 Radial Gradients in Stellar Population Parameters
Carbon (C) and nitrogen (N) are not considered $\alpha$ elements, but they follow the same qualitative trend as Mg. Radial gradients (as a function of $\sigma$) for C and N are shown for all the $\sigma$ bins in Figures 18 and 19 respectively. One notable feature in both of these figures is that for massive ETGs (log($\sigma$)$>$2.0) the radial gradients in C and N seem to have flattened with $\sigma$ on average, although the gradient amplitude is very small for these elements. Parikh et al. (2021) reported a very flat gradient for both of these elements for almost all of their mass bins. Feldmeier-Krause et al. (2021) reported a very flat gradient for C but slightly negative gradient for N. Within the errors, our results agree with these studies. Additionally, Feldmeier-Krause et al. (2021) found no significant correlation between radial gradients in C and N with [Fe/H], which we also did not find (figure not shown in this article).
@@
The strongest gradient values are obtained for sodium. Fig. 20 shows that the negative radial gradient in Na increases monotonically with $\sigma$. The median negative gradient of around 0.10 [Na/Fe]/log(R/Re) in this work is similar to a number of previously reported values La Barbera et al. (2017); Parikh et al. (2021). Feldmeier-Krause et al. (2021) reports strong negative radial gradient in [Na/R] for individual ETGs, but their [Na/R] values are higher than what we report in this work. Given the overall similarity between Mg and Na trends among ETGs, the $\sim 2$ times stronger Na gradients within ETGs are puzzling.
@@
Magnesium (Mg) is considered an $\alpha$-element that is predominantly produced inside massive stars via fusion of helium nuclei. In this work, Mg shows negative radial gradient for all $\sigma$ bins (qualitatively in Fig17 and quantitatively in Fig. 21). Early work by Davies et al. (1993) showed that the radial gradient in the Mg2 index has a value of around -0.06 which can be translated roughly to a radial gradient in [Mg/R]. Our highest $\sigma$ bin has a mean radial gradient value of around -0.02 [Mg/R]/log(R/Re) and gradually decreases to a flat gradient for the lowest $\sigma$ bin. Many previous studies reported a radial gradient value of near -0.06 (Rawle et al., 2008; Ferreras et al., 2019; Feldmeier-Krause et al., 2021), whereas Mehlert et al. (2003) and Parikh et al. (2021) report zero radial gradient for [Mg/R]. In the average ETG, if there is such a thing, the Mg abundance is global; independent of radial distances from the center of the galaxy.
@@ 5.3 Scatter in Astrophysical Parameters and their Radial Gradients
Once observational effects are subtracted, the scatter of a derived quantity within a group of galaxies reveals the extent of variation within the population allowed by natural evolutionary pathways. With over 2000 galaxies, sample size should not be a concern, and the extent of natural variation should be well-measured. In this section we show results of our calculated intrinsic scatter ($\delta_{i}$) and how they correlate with $\sigma$. In general, we find that observational errors are small enough that almost all of the scatter in previous figures comes from the intrinsic scatter.
@@
The best way to estimate $\delta_{o}$ is via repeated observations of the same target, which MaNGA does not provide. However, MaNGA does calculate the uncertainty per spectral pixel, and this is readily translated to uncertainty per measured index value. We assume that spectral response ripples over spectral scales of 10Å to 100Å, which can be a significant error source for some spectrographs, are not a significant source of error for MaNGA. Our method to estimate $\delta_{o}$ is to randomly vary each index measurement within its Gaussian probability distribution, and then recalculate compfit solutions. The as-measured solution was taken as the mean, then 50 variations for each annulus were generated. This provided us with a proxy for standard deviation for each measurement. To obtain an equivalent error for the radial gradient measurements, we used the error from the covariance matrix that was generated while calculating the radial gradients.
@@
In Fig. 23 we show the calculated $\delta_{i}$ for the age and radial gradient in age. We see a constant decreasing trend for $\delta_{i}$ in age with $\sigma$ and $\delta_{i}$ for the highest $\sigma$ bin (0.05 dex) is four times less than for the lowest $\sigma$ bin (0.2 dex). The $\delta_{i}$ in radial gradient trends the opposite way, with the flat gradients in low-$\sigma$ galaxies showing little scatter, while the almost-flat age gradients in high-$\sigma$ galaxies show large intrinsic scatter that straddles zero.
The slightly contradictory observation that the most age-homogeneous galaxies also have the most astrophysical spread in age gradient actually fits well with the scenario where the more massive galaxies are formed by large number of merger events compared to smaller galaxies which are considered to have gone through less number of merger events. Faber et al. (2007) showed that number of red galaxies (ETGs) have been significantly rising from $z\sim$1 to the present day ($z$=0). In order to explain this scenario, the authors posit that migration of galaxies to the red sequence is caused by both quenching and merger events. We also support a more homogeneous age distribution for galaxies with high $\sigma$. In a merger-dominated scenario, the placement of youthful stars post-merger (nearer to the center or nearer to the edges) seems dependent on gas content, kinematic initial conditions, and the ages of the merger components. That is, we see no contradiction that increased merging causes larger scatter in age gradient.
@@
Some galaxy formation scenarios posit that a gas collapse forms strong abundance gradients in galaxy subunits, many of which later merge and therefore lessen the gradients. These scenarios predict the weakest gradients should be seen in high-$\sigma$ galaxies. If gas is involved in the mergers, high-$\sigma$ galaxies should also show the highest abundances and youngest ages. None of these predictions are consistent with our results. A comparison with IllustrisTNG hierarchical simulations, on the other hand, reproduce the age structure we find, show the abundance slope changes at log($\sigma$$\approx$2.0 that we observe, and exhibit flat gradients similar to those we derive, although the abundances predicted by IllustrisTNG are significantly higher than our observations overall. The Illustris comparison is discussed below in section 5.4.
@@
Fig. 25 shows the $\delta_{i}$ values in C, N, Na, and Mg abundances. The intrinsic scatter in C abundance is the least, with Mg a close second. Apart from the first $\sigma$ bin, N abundance has nearly constant $\delta_{i}$ of 0.1 dex for all the other $\sigma$ bins. The high value of $\delta_{i}$ in N for smaller galaxies (log($\sigma$)$<$2.0) might shed light on the time-scale of chemical enrichment in them. If N from intermediate-mass stars is important, a variation in N abundance implies a variation in formation timescale. It is plausible that these variations would be attenuated in galaxies with richer merger histories. The $\delta_{i}$ for Na abundance is $\sim$0.14 dex for the first three $\sigma$ bins and decreasing marginally to 0.12 dex for the last two high $\sigma$ bins. This behavior of Na is not yet fully understood, but additional absorption of Na in the interstellar medium (ISM) could increase inferred [Na/R] in some galaxies and increase scatter.
@@
We show the $\delta_{i}$ values in radial gradients of C, N, Na, and Mg abundances in Fig. 26. The most interesting result is the increasing trend of $\delta_{i}$ from the second $\sigma$ bin onwards till the penultimate $\sigma$ bin, reaching a value of almost a factor of two higher compared to the second $\sigma$ bin. Surprisingly, the $\delta_{i}$ in all the elements for the highest $\sigma$ bin drops to a lower value. Radial gradients in C and Mg have the lowest values in $\delta_{i}$ followed by N and then Na. This similar trend is also seen in Fig. 25. It is noteworthy that radial gradients in Na have significantly higher $\delta_{i}$ compared to other elements considered in this work reaching as high as 0.22 dex/log(R/Re) in the fourth $\sigma$ bin. In conjunction with the Fig. 20, it can be stated that there is a significant and as-yet-unexplained anomaly regarding Na, which has not been sufficiently addressed in the existing body of literature. ISM absorption may play a small role, but the overall coherence of Na with Mg and other elements argues that ISM absorption is the exception rather than the rule.
@@
[1.6, 2] 0.20 0.15 0.06 0.15 0.13 0.08 0.10 0.15 0.05 0.14 0.15 0.06
[2, 2.2] 0.18 0.11 0.06 0.12 0.14 0.06 0.12 0.14 0.04 0.07 0.10 0.04
[2.2, 2.34] 0.15 0.09 0.05 0.10 0.13 0.06 0.20 0.14 0.06 0.10 0.15 0.05
[2.34, 2.45] 0.10 0.11 0.07 0.12 0.12 0.07 0.35 0.19 0.08 0.15 0.22 0.09
[2.45, 2.53] 0.07 0.10 0.04 0.09 0.12 0.06 0.45 0.22 0.06 0.10 0.15 0.08
@@
This observed trend of decreasing metallicity with increasing age is a result of chemical evolution of the galaxies and their star formation history. Perhaps, elliptical galaxy precursors formed most of their stars early in the Universe’s history through rapid, intense starbursts (Partridge & Peebles, 1967; De Lucia et al., 2006). At that early stage of Universe’s history, the relatively metal-poor gas formed metal poor stars. Elliptical galaxies are often the product of major mergers of smaller galaxies (Toomre & Toomre, 1972; Taranu et al., 2015). These mergers can trigger intense star formation early in a galaxy’s evolution but can also lead to the expulsion of gas through galactic winds, particularly if supernovae and active galactic nuclei (AGN) are involved (Loewenstein & Mathews, 1987; Forman et al., 1990). Thus the decrease in metallicity with increasing age in elliptical galaxies might be explained as a result of rapid early star formation, feedback processes that expel metal-rich gas, and the cessation of star formation early in the galaxy’s history.
5.4 Comparison with IllustrisTNG
The Next Generation Illustris (IllustrisTNG) project is a series of magnetohydrodynamical cosmological simulations designed to investigate galaxy formation and evolution within the framework of the $\Lambda$ Cold Dark Matter ($\Lambda$CDM) cosmological model (Marinacci et al., 2018; Naiman et al., 2018; Nelson et al., 2018; Pillepich et al., 2018; Springel et al., 2018). It builds and improves on the scientific insights and achievements of the original Illustris simulation (Vogelsberger et al., 2014b; Genel et al., 2014). Nine individual elements – hydrogen (H), helium (He), carbon (C), nitrogen (N), oxygen (O), neon (Ne), magnesium (Mg), silicon (Si), and iron (Fe) – were explicitly tracked throughout the IllustrisTNG simulation (Naiman et al., 2018). In this work, we compare our results with those of publicly available IllustrisTNG data (Nelson et al., 2019). Only a couple of studies have so far compared the results of the IllustrisTNG simulations with the observations (Naiman et al., 2018; Pulsoni et al., 2020). We sought to eliminate this scarcity as an additional task in this work.
We compare our results with that of IllustrisTNG in two different ways: (1) grouped by $\sigma$ following the binning scheme outlined in the Table 2 and plotted versus $R/R_{e}$, and (2) ungrouped versus $\sigma$ using the 3rd annulus for the MaNGA galaxies. We provide all the comparison plots in Appendix A.
Mean age comparison. Fig. 28 shows the comparison for IllustrisTNG and MaNGA ages. In the left-hand panel of the plot we see that the IllustrisTNG galaxies (solid lines) in all $\sigma$ bins show a very flat radial trend in age and it is very well matches our calculations (stars with errorbars) of age radial gradients within each $\sigma$ bin. The right-hand panel shows the age of each individual IllustrisTNG galaxy (in red cross) plotted against their $\sigma$, overlaid on our findings (in blue star) for the age at 3rd annulus for each MaNGA galaxy. We again see striking similarity between the IllustrisTNG results and our work. Although the scatter in IllustrisTNG measurements is less across all $\sigma$ values compared to MaNGA, it is interesting to note that scatter decreases with increasing $\sigma$ for IllustrisTNG simulation as well.
Mean [Fe/H] comparison. Two panels in Fig. 29 shows comparison between IllustrisTNG and MaNGA results for [Fe/H]. Except for a zero point shift of about a factor of two (IllustrisTNG being more metal rich), the morphologies of the trends are remarkably similar. The spatial gradients are quantitatively similar, and even show the steepening with increasing $\sigma$ (left-hand panel). IllustrisTNG results also show a decreasing trend in [Fe/H] for galaxies with log($\sigma$)$>$2.0 (right-hand panel). Our MaNGA sample might place the inflection point at log($\sigma$)$>$2.1, instead, but the existence of the inflection is a remarkable success for IllustrisTNG. Because compfit will not return values of [Fe/H] higher than 0.4, we cannot comment on the small number of IllustrisTNG galaxies that scatter to higher [Fe/H] values. We also observe a $\sim$1fraction of galaxies that scatter to low abundance. These are not reflected the theoretical simulation.
Mean [X/Fe] comparison. While considering the individual elements (Fig. 30 for C, Fig. 31 for N, and Fig. 32 for Mg), the most striking aspect is the zeropoint offsets. To reiterate our methods, we include magnesium amplification for the first time, so our abundance results should approach actual [X/Fe] values for the first time in the literature. IllustrisTNG, on the other hand, will have adjusted their assumptions of stellar element yields to follow earlier estimates. Therefore, these zeropoint offsets should be no surprise. Parenthetically, Naiman et al. (2018) also reported an overabundance of magnesium for Milky Way-like galaxies in the IllustrisTNG simulation. Looking past that, the shapes of the data clouds are qualitatively similar. The abundance inflection at log $(\sigma)=2$ is seen in all elements in both theory and observation.
The inflection at log $(\sigma)\approx 2$. The fact that IllustrisTNG simulations qualitatively reproduce the chemical “phase change” at log $(\sigma)\approx 2$ is a success of the theory. To our knowledge, no author has pointed out this particular success before. Low-$\sigma$ galaxies are effective at Fe enrichment, but high-$\sigma$ galaxies favor light metal enrichment to such a degree that [Fe/H] falls with increasing $\sigma$. By itself, this implies that high=$\sigma$ galaxies cannot simply be dry mergers of $\sim$100 km s-1 subunits: the inherited abundance pattern does not persist. It is also hard to imagine a monolithic collapse scenario in which galaxies with shallower potential wells retain more iron, since supernova winds should be relatively more effective in low-$\sigma$ galaxies. But IllustrisTNG demonstrates that the hierarchical merger picture can thread the needle between extremes and reproduce the age-$\sigma$ trend along with the chemical inflection at log $(\sigma)\approx 2$.
@@ 5.4 Comparison with IllustrisTNG
Nitrogen. The large scatter in N, especially compared to IllustrisTNG, is remarkable. Because the scatter in [N/Fe] is astrophysical, it is clear that much information is being carried in this tracer element regarding the chemical histories of galaxies. Nitrogen of some quantity is produced instantly in supernovae, at young ages in WR stars and other massive stars, at intermediate ages via CNO processed material, and even in old red giants by mass loss after first dredge-up. Due to the unsolved status of mass loss in stars, the balancing of yields from these sources is not predictable from first principles. However, our results here hint at the importance of CNO processed material. Since increased N comes at the expense of C in the CNO cycle, one might expect an anticorrelation between the two abundances, other things being equal. In fact, we nearly observe this because the correlation between [N/R] and [C/R] gradients is the weakest. In the annulus 3 plots, [N/R] correlates least tightly with [C/R] and [Na/R]. The possibility that we see the effects of CNO processing in these data is real.
@@
ADF comparison. We compared the Abundance Distribution Function (ADF) used in this study with six halos taken from IllustrisTNG, as shown in Fig. 33. The analytical ADF formula was taken from Worthey et al. (2005) and Tang et al. (2014), and we used the ‘normal’ width ADF to derive peak [Fe/H] from MaNGA galaxies. The ‘normal’ ADF has FWHM width of 0.65 dex. The six IllustrisTNG ADFs average to FWHM = 0.67 dex, and the overall shapes are similar. This shows the consistency between our assumptions, which are based on observations in the local universe, and the IllustrisTNG calculations. Our composite models should be free of bias as compared to IllustrisTNG, at least along the dimension of abundance-compositeness. On the other hand, the overall chemical evolution assumptions in IllustrisTNG give ADFs that resemble observations, except for a small parcel of relic stars at very low metallicity (Population III stars, not illustrated in Fig. 33) that are so far unseen in the local universe, and therefore may not exist.

### v2

@@ 5.2 Radial Gradients in Age and Abundances
Although carbon (C) and nitrogen (N) are not classified as $\alpha$ elements, their radial gradients mirror the qualitative trend of magnesium (Mg). Figs. 19 and 20 depict the radial gradients of C and N, respectively, across all $\sigma$ bins. Notably, for massive early-type galaxies (ETGs) with log $\sigma>2.0$, the radial gradients of C and N exhibit flattening with respect to $\sigma$, albeit with minimal gradient amplitude. This aligns with Parikh et al. (2021), who reported nearly flat gradients for both elements across most mass bins, and Feldmeier-Krause et al. (2021), who noted a flat gradient for C and a slightly negative gradient for N. Within uncertainties, our findings are consistent with these studies. Furthermore, like Feldmeier-Krause et al. (2021), we find no significant correlation between C and N radial gradients and [Fe/H] (figure not included).
@@
The strongest gradient values are obtained for sodium. Fig. 21 shows that the negative radial gradient in Na increases monotonically with $\sigma$. The median negative gradient of around 0.10 dex per decade ([Na/R]/log R/Re) in this work is similar to a number of previously reported values La Barbera et al. (2017); Parikh et al. (2021). Feldmeier-Krause et al. (2021) reports strong negative radial gradient in [Na/R] for individual ETGs, but their [Na/R] values are higher than what we report in this work. Given the overall similarity between Mg and Na trends among ETGs, the $\sim 2$ times stronger Na gradients within ETGs are puzzling.
Let us briefly consider absorption of Na in the cold interstellar material (ISM). The Na D index measures a pair of transitions whose lower state is the ground state, i.e., a resonant feature in the ISM. Absorption along lines of sight out of the Milky Way plane is on the order of tenths of Å (Bica & Alloin, 1986) compared with observed ETG Na D strengths of $\sim$3 Å. While more massive ETGs tend to harbor more cool ISM (Welch et al., 2010; Werner et al., 2014) in the outskirts, the amount of Na would have to be considerable. Also, more absorption in the outskirts would lead to weaker, not stronger, negative radial gradients. All things considered, we judge that cold interstellar Na is not a plausible explanation for the [Na/R] gradient trends in ETGs.
@@
Magnesium (Mg) is considered an $\alpha$-element that is predominantly produced inside massive stars via fusion of helium nuclei. In this work, Mg shows negative radial gradient for all $\sigma$ bins (qualitatively in Fig18 and quantitatively in Fig. 22). Early work by Davies et al. (1993) showed that the radial gradient in the Mg2 index has a value of around -0.06 which can be translated roughly to similar radial gradient values in [Mg/R]. Our highest $\sigma$ bin has a mean radial gradient value of around -0.02 dex per decade ([Mg/R]/log R/Re) and gradually decreases to a flat gradient for the lowest $\sigma$ bin. Many previous studies reported a radial gradient value of near -0.06 (Rawle et al., 2008; Ferreras et al., 2019; Feldmeier-Krause et al., 2021), whereas Mehlert et al. (2003) and Parikh et al. (2021) report zero radial gradient for [Mg/R]. In the average ETG, if there is such a thing, the Mg abundance is global; independent of radial distances from the center of the galaxy.
The nearly flat radial gradients of C, N, and Mg in ETGs support a hierarchical assembly scenario, where repeated mergers and accretion events mix stars and gas and spatially homogenize the stellar populations. This contrasts with the steep gradients expected from a simple monolithic collapse. The shallow light element gradients favor mergers as agents to homongenize ETGs and favor their formation as extended, multi-phase events rather than rapid, single-collapse systems.
@@ 5.3 Scatter in Astrophysical Parameters and their Radial Gradients
Once observational effects are subtracted, the scatter of a derived quantity within a group of galaxies reveals the extent of variation within the population allowed by natural evolutionary pathways. With over 2000 galaxies, our sample size should not be a concern, and the extent of natural variation should be well-measured. In this section, we show results of our calculated intrinsic scatter ($\delta_{i}$) and how they correlate with $\sigma$. In general, we find that observational errors are small enough that almost all of the scatter in previous figures comes from the intrinsic scatter.
@@
Optimal estimation of $\delta_{o}$ relies on repeated observations of the same target, which MaNGA does not provide. However, MaNGA yields per-spectral-pixel uncertainty, readily convertible to uncertainty per measured index value. We assume that spectral response ripples over 10Å to 100Å scales, a potential error source for some spectrographs, are negligible for MaNGA. To estimate $\delta_{o}$, we randomly perturbed each index measurement within its Gaussian probability distribution and recalculated compfit solutions. The mean was derived from the as-measured solution, with 50 variations (MC realizations)) generated per annulus to approximate the standard deviation for each measurement. For radial gradient measurement errors, we utilized the covariance matrix error obtained during gradient calculations.
@@
In Fig. 25 we show the calculated $\delta_{i}$ for the age and radial gradient in age. The figure shows a slightly decreasing trend with $\sigma$ for age scatter whereas the opposite is observed for age gradient scatter. The flat gradients in low-$\sigma$ galaxies show little scatter ($\sim$0.09 dex/log(R/Re)), while the almost-flat age gradients in high-$\sigma$ galaxies show slightly larger intrinsic scatter ($\sim$0.15 dex/log(R/Re)). Sliced the other way, low-$\sigma$ galaxies are well-mixed radially, but globally show substantial age scatter while high-$\sigma$ galaxies have small but measurable age gradients, both negative and positive, but globally are older and more homogeneous in their senescence.
The slightly contradictory observation that the most age-homogeneous galaxies also have one of the highest astrophysical spreads in age gradient actually fits well with the scenario where the more massive galaxies are formed by large number of merger events compared to smaller galaxies which are considered to have gone through fewer merger events. Faber et al. (2007) showed that the number of red galaxies (ETGs) has increased from $z\sim$1 to the present day (zz=0). In order to explain this scenario, the authors posit that migration of galaxies to the red sequence is caused by both quenching and merger events. We also support a more homogeneous age distribution for galaxies with high $\sigma$. In a merger-dominated scenario, the placement of young stars post-merger (nearer to the center or nearer to the edges) might depend on gas content, kinematic initial conditions, and the ages of the merger components. That is, it is plausible that increased merging causes larger scatter in age radial gradients.
@@
[…]
@@
Fig. 27 shows the $\delta_{i}$ values in C, N, Na, and Mg abundances. The intrinsic scatter in C abundance is the least, with Mg a close second. Apart from the first $\sigma$ bin, N abundance has nearly constant $\delta_{i}$ of 0.1 dex for all the other $\sigma$ bins. The high value of $\delta_{i}$ in N for smaller galaxies (log($\sigma$)<<2.0) might shed light on the time-scale of chemical enrichment in them. If N from intermediate-mass stars is important, a variation in N abundance implies a variation in formation timescale. It is plausible that these variations would be attenuated in galaxies with richer merger histories. The $\delta_{i}$ for the Na abundance declines from $\sim$0.14 dex for small galaxies to $\sim$0.12 dex for large ones. Unique to Na is that additional absorption of Na in the interstellar medium (ISM) could increase inferred [Na/R] in some galaxies and increase scatter.
@@
The $\delta_{i}$ values for radial gradients in C, N, Na, and Mg abundances appear in Fig. 28. The most interesting result is the increasing trend of $\delta_{i}$ from the second $\sigma$ bin onwards till the penultimate $\sigma$ bin, reaching a value of almost a factor of two higher compared to the second $\sigma$ bin. Surprisingly, the $\delta_{i}$ in all the elements for the highest $\sigma$ bin drops to a lower value. Radial gradients in C and Mg have the lowest values in $\delta_{i}$ followed by N and then Na. This similar trend is also seen in Fig. 27. It is noteworthy that radial gradients in Na have significantly higher $\delta_{i}$ compared to other elements considered in this work reaching as high as 0.22 dex/log(R/Re) in the fourth $\sigma$ bin. In conjunction with the Fig. 21, it can be stated that there is a significant and as-yet-unexplained anomaly regarding Na, which has not been sufficiently addressed in the existing body of literature. ISM absorption may play a small role, but the overall coherence of Na with Mg and other elements argues that ISM absorption is the exception rather than the rule.
@@
[1.6, 2] 0.13 0.17 0.06 0.15 0.14 0.09 0.09 0.15 0.07 0.15 0.16 0.09
[2, 2.2] 0.11 0.13 0.06 0.11 0.15 0.07 0.07 0.18 0.05 0.08 0.12 0.05
[2.2, 2.34] 0.11 0.13 0.06 0.10 0.13 0.07 0.09 0.19 0.06 0.11 0.16 0.06
[2.34, 2.45] 0.14 0.13 0.07 0.11 0.12 0.07 0.17 0.23 0.09 0.16 0.21 0.17
[2.45, 2.53] 0.07 0.12 0.06 0.09 0.12 0.06 0.14 0.25 0.07 0.14 0.17 0.08
We now provide brief remarks on the age-[Fe/H] relation. Fig. 29 shows the distribution of age with [Fe/H] for each annulus grouped by [Fe/H] (top panel) and for the 3rd annulus of each individual galaxy (bottom panel). The $\delta_{i}$ in age and [Fe/H] is sufficient to explain the errorbars in top panel of Fig. 29. We see a decreasing trend of age with [Fe/H] (at least for the last three [Fe/H] bins). The trendline shows that the age of ETGs decrease by as much as 0.57 dex (equivalent to 3.7 Gyr) for an increase in [Fe/H] by 1 dex. One of the first works on characterizing age-[Fe/H] relation was done by Worthey et al. (1995). They showed that E/S0 galaxies follow the direction of null spectral change characterized by $\DeltalogZZ/Δ\Deltalog$ age = -2/3. The slope of -0.57 dex/dex that we derived in log age - [Fe/H] plane (bottom panel of Fig. 29) is not too far away from the numerical value of -2/3 ($\approx$-0.67). A very similar result was reported by other studies that looked into elliptical galaxies in Fornax cluster (Kuntschner, 2000) and Coma cluster (Jørgensen, 1999; Poggianti et al., 2001). Gallazzi et al. (2005) analyzed almost 26,000 ETGs from SDSS-DR2 and showed that (their Fig. 12) the age-[Fe/H] relation have an anti-correlation for all of their mass bins. Galaxies in different environments also show the trend from old and metal-poor to young and metal-rich in Coma supercluster (Tiwari et al., 2020). Although any trend that flirts with being parallel to the age-metallicity degeneracy is fraught with modeling systematics, the variety of models employed to find the same result is now significant, so perhaps we should begin to accept it as a genuine constraint on galaxy evolution codes like TNG.
This observed trend of decreasing metallicity with increasing age may be a straightforward result of chemical evolution. Elliptical galaxy precursors are thought to form most of their stars early in the Universe’s history through rapid, intense starbursts (Partridge & Peebles, 1967; De Lucia et al., 2006). At that early stage of Universe’s history, the relatively metal-poor gas formed metal poor stars. Elliptical galaxies are often the product of major mergers of smaller galaxies (Toomre & Toomre, 1972; Taranu et al., 2015). These mergers can trigger intense star formation early in a galaxy’s evolution but can also lead to the expulsion of gas through galactic winds, particularly if supernovae and active galactic nuclei (AGN) are involved (Loewenstein & Mathews, 1987; Forman et al., 1990). However, if the gas does not escape altogether, it will only become more chemically enriched with time, form stars, and create an age-metallicity relation. The question then turns toward matching the amplitude of what is observed with galaxy evolution models.
5.4 Comparison with TNG
The Next Generation Illustris (TNG) project is a series of magnetohydrodynamical cosmological simulations designed to investigate galaxy formation and evolution within the framework of the $\Lambda$ Cold Dark Matter ($\Lambda$ -CDM) cosmological model (Marinacci et al., 2018; Naiman et al., 2018; Nelson et al., 2018; Pillepich et al., 2018; Springel et al., 2018). It builds and improves on the scientific insights and achievements of the original TNG simulation (Vogelsberger et al., 2014b; Genel et al., 2014). Nine individual elements – hydrogen (H), helium (He), carbon (C), nitrogen (N), oxygen (O), neon (Ne), magnesium (Mg), silicon (Si), and iron (Fe) – were explicitly tracked throughout the TNG simulation (Naiman et al., 2018). In this work, we compare our results with those of publicly available TNG data (Nelson et al., 2019). Note that we use the TNG-100 simulation results for comparison with our observations. Only a couple of studies have so far compared the results of the TNG simulations with the observations (Naiman et al., 2018; Pulsoni et al., 2020).
We compare our results with that of TNG in two different ways: (1) grouped by $\sigma$ following the binning scheme outlined in the Table 2 and plotted versus $R/R_{e}$, and (2) ungrouped versus $\sigma$ using the 3rd annulus for the MaNGA galaxies. We provide all the comparison plots in Appendix A.
Mean age comparison. Fig. 30 shows the comparison for TNG and MaNGA ages. In the left-hand panel of the plot we see that the TNG galaxies (solid lines) in all $\sigma$ bins show a flat radial trend in age similar to our calculations (stars with errorbars) of age radial gradients within each $\sigma$ bin. The right-hand panel shows the age of each individual TNG galaxy (in red cross) plotted against their $\sigma$, overlaid on our findings (in blue star) for the age at 3rd annulus for each MaNGA galaxy. The TNG simulations show an old age “ceiling” that the MaNGA observations do not mimic except for high-$\sigma$ galaxies. The scatter in TNG measurements is less across all $\sigma$ values compared to MaNGA. However, scatter decreases with increasing $\sigma$ for TNG simulation as well.
Mean [Fe/H] comparison. Two panels in Fig. 31 shows comparison between TNG and MaNGA results for [Fe/H]. The morphologies of the trends are remarkably similar. The spatial gradients are quantitatively similar (left-hand panel). The TNG results also show a decreasing trend in [Fe/H] for galaxies with log($\sigma$)>>2.0 (right-hand panel). The existence of this inflection is a remarkable success for TNG. Because compfit will not return values of [Fe/H] higher than 0.4, we cannot comment on the small number of TNG galaxies that scatter to higher [Fe/H] values. We also observe a $\sim$1% fraction of galaxies that scatter to low abundance. These are not reflected the theoretical simulation.
Mean [X/Fe] comparison. While considering the individual elements (Fig. 32 for C, Fig. 33 for N, and Fig. 34 for Mg), we see zeropoint offsets along with much less scatter in TNG simulations compared to MaNGA results. We include magnesium amplification for the first time, so our abundance results should approach actual [X/Fe] values for the first time in the literature. TNG, on the other hand, will have adjusted their assumptions of stellar element yields to follow earlier estimates. Therefore, the existence of zeropoint offsets should be no surprise. The offset is strongest for magnesium. Parenthetically, Naiman et al. (2018) also reported an overabundance of magnesium for Milky Way-like galaxies in the TNG simulation. A hint of abundance inflection at log $(\sigma)=2$ is seen in all elements in both theory and observation.
@@ 5.4 Comparison with TNG
Nitrogen. The large scatter in N, especially compared to TNG, is remarkable. Because the scatter in [N/Fe] is astrophysical, it is clear that much information is being carried in this tracer element regarding the chemical histories of galaxies. Nitrogen of some quantity is produced (1) instantly in supernovae, (2) at young ages in WR stars and other massive stars, (3) at intermediate ages via CNO processed material, (4) and in old red giants by mass loss after first dredge-up. Due to the unsolved status of mass loss in stars, the balancing of yields from these sources is not predictable from first principles. However, our results here hint at the importance of CNO processed material. Since increased N comes at the expense of C in the CNO cycle, one might expect an anticorrelation between the two abundances, other things being equal. In fact, we might observe this because the correlation between [N/R] and [C/R] gradients is the weakest of any element gradient combination. In the annulus 3 plots, [N/R] correlates least tightly with [C/R] and [Na/R]. The possibility that we see the effects of CNO processing in these data is real.
@@
ADF comparison. We compared the Abundance Distribution Function (ADF) used in this study with six halos taken from TNG, as shown in Fig. 35. The analytical ADF formula was taken from Worthey et al. (2005) and Tang et al. (2014), and we used the ‘normal’ width ADF to derive peak [Fe/H] from MaNGA galaxies. The ‘normal’ ADF has FWHM width of 0.65 dex. The six TNG ADFs average to FWHM = 0.67 dex, and the overall shapes are similar. Our ADF assumption, which is based on observations in the local universe, seem consistent with TNG calculations. Therefore, our composite models should be free of bias as compared to TNG, at least along the dimension of abundance-compositeness. The overall chemical evolution assumptions in TNG give ADFs that resemble observations, except for a small parcel of relic stars at very low metallicity (Population III stars, not illustrated in Fig. 35) that are so far unseen in the local universe and therefore may not exist.